# Head-wise Shareable Attention for Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) suffer from huge number of parameters, which restricts their deployment on edge devices. Weight sharing is one promising solution that encourages weight reuse, effectively reducing memory usage with less performance drop. However, current weight sharing techniques primarily focus on small-scale models like BERT and employ coarse-grained sharing rules, e.g., layer-wise. This becomes limiting given the prevalence of LLMs and sharing an entire layer or block obviously diminishes the flexibility of weight sharing. In this paper, we present a perspective on head-wise shareable attention for large language models. We further propose two memory-efficient methods that share parameters across attention heads, with a specific focus on LLMs. Both of them use the same dynamic strategy to select the shared weight matrices. The first method directly reuses the pre-trained weights without retraining, denoted as DirectShare. The second method first post-trains with constraint on weight matrix similarity and then shares, denoted as PostShare. Experimental results reveal our head-wise shared models still maintain satisfactory capabilities, demonstrating the feasibility of fine-grained weight sharing applied to LLMs.
@@ 1 Introduction
Large Language Models (LLMs) have achieved breakthrough performance in a variety of natural language processing tasks  (Wei et al., 2022; Bubeck et al., 2023; Zhao et al., 2023). However, such remarkable capability typically comes at the cost of a substantial increase in the model size  (Kaplan et al., 2020). Thus, LLMs with billions of parameters  (Brown et al., 2020; Touvron et al., 2023) are more resource-hungry despite a wide margin of superiority over small-scale models (Devlin et al., 2018; Liu et al., 2019). This can also pose challenges for deployment on low-capability devices due to limited storage and GPU memory.
To address the high memory requirements of models, weight sharing  (Takase and Kiyono, 2021; Liu et al., 2023) aims to reuse the same parameters to achieve memory- and storage-efficiency while preserving model performance. For small-scale models, e.g., BERT, it is known that several techniques  (Lan et al., 2019; Liu et al., 2023) are proposed to explore across-layer parameter sharing. While,  Zhang et al. (2022) demonstrate identical weights across different layers are the main cause of training instability and performance degradation. Moreover, the effective of similar techniques at the scale of LLMs remains uncertain.
Thus, we strive to solve this central question: Can we design fine-grained weight sharing strategy that can smoothly apply to large language models? For an effective memory-efficient weight sharing method tailored to LLMs, two key challenges must be tackled: a) the choice of shared modules whose weights are reused; b) the trade-off between reducing memory footprint and preserving diverse capabilities.
@@ 2.1 Memory-efficient Approaches for LLMs
With the growing size of language models, several memory-efficient techniques are proposed to solve. One line to reducing the memory footprint involves network compression, like quantization  (Bai et al., 2020; Tao et al., 2022), pruning  (Yang et al., 2022; Tao et al., 2023) and knowledge distillation (Wu et al., 2023; Tan et al., 2023). However, when applied to LLMs, many approaches have become infeasible  (Frantar and Alistarh, 2023). To recover accuracy, they require extensive post-training of the model  (Dettmers et al., 2023; Sun et al., 2023).
@@ 2.2 Weight Sharing
Weight sharing is a widely used technique  (Lan et al., 2019; Liu et al., 2023; Lv et al., 2023; Xu and McAuley, 2023) that aims to improve parameter efficiency and reduce inference memory footprint. Weight sharing enables model compression by eliminating redundant parameters and decouples computation and parameters by reusing the same parameters for multiple computations.
Task-oriented Weight Sharing. One of the prevalent tasks using weight sharing mechanisms is nerual machine translation (NMT). Tied Transformer  (Xia et al., 2019) considers model-level sharing and shares the weights of the encoder and decoder of an NMT model.  Dabre and Fujita (2019) proposes a method, which shares the weights across all Transformer layers and keeps performance in NMT. Besides,  Chi et al. (2021) bring the idea of ALBERT  (Lan et al., 2019) to the speech recognition task.
[…]
@@ 3.1 Attention Map Similarity: From Layer-wise to Head-wise
Prior researches (Xiao et al., 2019; Ying et al., 2021; Bhojanapalli et al., 2021) demonstrate the effectiveness of attention map reuse due to the high similarity of attention scores between different layers (especially for adjacency layers). Motivated by this, we delves into attention map similarity, specifically transitioning from layer-wise to head-wise analysis. To measure the evolution of the attention maps over layers and heads, we use the cosine similarity $\mathcal{S}_{cos}$. When $\mathcal{S}_{cos}$ equals one, it means that the attention maps are perfectly similar. Considering two specific self-attention layers, the cosine similarity is calculated as follows:
@@
We visualize the layer-wise and head-wise attention map similarity across three task-specific datasets: WMT14 (En-Fr) (Bojar et al., 2014), CommonsenceQA (Talmor et al., 2019) and WSC (Levesque et al., 2012). As shown in Fig. 1(a) and (b), the degrees of similarity in attention scores computed in different layers and heads present a certain level of consistency across different tasks. In addition, we find that the cosine similarity values for pairs with high similarity are higher among different heads compared to different layers. Specifically, the most similar self-attention layers reach a cosine similarity value of approximately 0.90, while in the case of head-wise comparisons, several pairs have a remarkable similarity of nearly 0.99.
One observation is that as the number of parameters increases, modules with high similarity exhibit variations, particularly in the fine-grained (e.g., head-wise) comparisons within large-scale pre-trained language models. Existing approaches employ "learning to share" techniques to dynamically adjust the sharing strategy (Xiao et al., 2019) or use a uniform sharing strategy but train the modified model from scratch (Ying et al., 2021; Shim et al., 2023). However, such strategies pay little attention on reusing attention map among heads and incur high computational costs for LLMs.
@@ 3.2 From Attention Map Similarity to Weight Matrix Similarity
Combined with the analysis of weight matrix similarity, we have made a key discovery: given a pre-trained LLM, by concatenating the weight matrix $W^{q}$ and $W^{k}$ for each head to measure the cosine similarity, the most similar weight matrix corresponds to the overlapping modules with highly similar attention maps observed across different datasets. As illustrated in Fig. 1(b) and (c), deep green circles mark the connection between attention map similarity and weight similarity.
@@ 4.1 Head-wise Weight Sharing Strategy
To investigate the strategy of weight sharing applied to all the above three weight matrices across heads for LLMs, we perform preliminary experiments in the choice of head-wise match functions Match(normal-$\cdot$,$\cdot$). For the match functions, inputs are the weight matrices of head $\bm{i,j}$ and outputs are called matching scores $\bm{m}$. The higher the score, the more likely it is to share parameters across the heads.
$\bm{m_{i,j}^{*}}=\textbf{Match(}W_{i}^{*},W_{j}^{*}\textbf{)},*\in\{q,k,v\}$ (2)
@@
Besides, we try another five match functions to compare: (1) Only $W^{q}_{i}$ used to measure the cosine similarity, i.e., $\bm{m_{i,j}^{*}}=\mathcal{S}_{cos}(W^{q}_{i},W^{q}_{j})$; (2) Only $W^{k}_{i}$ used to measure the cosine similarity, i.e., $\bm{m_{i,j}^{*}}=\mathcal{S}_{cos}(W^{k}_{i},W^{k}_{j})$; (3) Only $W^{v}_{i}$ used to measure the cosine similarity, i.e., $\bm{m_{i,j}^{*}}=\mathcal{S}_{cos}(W^{v}_{i},W^{v}_{j})$; (4) Concatenate all the three matrices and then calculate the cosine similarity, i.e., $\bm{m_{i,j}^{*}}=\mathcal{S}_{cos}(W^{q}_{i}||W^{k}_{i}||W^{v}_{i},W^{q}_{j}||W^{k}_{j}||W^{v}_{j})$; (5) Separately use $W^{q}_{i},W^{k}_{i},W^{v}_{i}$ to measure the cosine similarity and do weight sharing respectively, i.e., $\bm{m_{i,j}^{*}}=\mathcal{S}_{cos}(W^{*}_{i},W^{*}_{j})$ and again *$*\in\{q,k,v\}$.
@@ 4.2 DirectShare
In practice, we traverse all head pairs to compute matching scores on Equation 3 and for each head, select the one with the highest score to match. When $M$ candidate head pairs prepared, we select the top-$N$ pairs in descending order according to the desired sharing ratio. Finally, we can share the weight matrices together between each selected attention head pairs. A detailed algorithm for our DirectShare is presented in Algorithm 1.
Input: Sharing Ratio $\alpha$, Weight Matrices of the MHA $\left\{W^{*}\right\},*\in\left\{q,k,v\right\}$
Output: The LLM after weight sharing
1 for $layer_{i}\leftarrow 2$ to $l$ do
2      for $i\leftarrow 1$ to $head\_num$ do
3           $match\_index$leftarrow -$1$
4           $match\_score$leftarrow -$1$
@@
6                for $j\leftarrow 1$ to $head\_num$ do
7                     Compute $\bm{m_{i,j}^{*}}$ using Eq. 3;
8                     if $\bm{m_{i,j}^{*}}>match\_score$ then
9                          $match\_score\leftarrow\bm{m_{i,j}^{*}}$
10                          $match\_index\leftarrow(layer_{j},j)$
11
12
13          Get one candidate head pair $<i,match\_index>$ for sharing;
@@
15
16Sort the matching scores in descending order;
17 Select the top-$N$ pairs according to $\alpha$;
Share the weight matrices together between each selected attention head pairs.
@@ 4.3 PostShare
$\mathcal{L}_{w}=\frac{1}{N}\sum_{n=1}^{N}\left(\sum_{*\in\{q,k,v\}}\left\|W_{n,i}^{*}-W_{n,j}^{*}\right\|_{2}\right)$ (4)
where $N$ is the number of selected attention head pairs <$i$,$j$> for sharing. With this regularization weight loss, the proposed PostShare learn model weights $W$ by minimizing the following combined loss function:
$\min_{W}\mathcal{L}_{post-training}+\gamma\times\mathcal{L}_{w}$ (5)
@@ 5.1 Experimental Settings
Backbone Models. We evaluate DirectShare and PostShare on two open-source LLMs: Llama 2 (Touvron et al., 2023) and Baichuan 2 (Baichuan, 2023) with 7B and 13B parameters. In PostShare, we use English Wikipedia (Foundation, ) to post-train the backbone models for weight sharing.
@@
Baselines. Since existing weight sharing techniques do not support LLMs, we compare DirectShare against Magnitude Pruning (Zhu and Gupta, 2017) and LLM-Pruner (Ma et al., 2023), two influential works for model pruning. Certainly, they are not directly comparable. To ensure fairness in the experiments, both of them only prune the multi-head attention module and thus compare when the same number of parameters is reduced. See Appendix A for additional information.
@@ 5.2.1 Evaluation on DirectShare
Logical and Common Sense Reasoning. In the domain of reasoning, we consider two Chinese natural language inference benchmarks and three English benchmarks: CMNLI (Xu et al., 2020), OCNLI (Hu et al., 2020), along with AX-b, AX-g and RTE from SuperGLUE (Wang et al., 2020).
In Table 1, we show the results on the above five tasks of memory-efficient Llama 2 models. The corresponding results for Baichuan 2 models can be found in Appendix B.1. When applying a 30% parameter sharing to Llama 2-7B, our DirectShare can still maintain an average performance of 99.51% across the five benchmarks, compared to the base model. With the same setting, the shared Llama 2-13B retains 99.21% performance. This suggests our finding of head-wise shareable attention for LLMs indeed can work without significant performance degradation in reasoning tasks.
@@
Ratio Method CMNLI OCNLI AX-b AX-g RTE
0% Llama 2-7B 32.98 33.12 53.53 55.34 49.82
10% Magnitude 32.99 30.63 56.70 49.44 47.29
LLM-Pruner 32.99 33.75 57.61 50.00 48.38
DirectShare 33.00 32.50 54.17 51.97 50.90
30% Magnitude 33.16 35.00 54.71 50.56 46.93
LLM-Pruner 32.99 31.25 56.34 52.53 48.74
DirectShare 33.33 32.50 57.07 51.69 49.10
0% Llama 2-13B 32.99 35.00 58.81 50.56 47.29
10% Magnitude 32.82 33.12 51.99 50.56 48.38
LLM-Pruner 32.99 36.25 58.70 50.00 46.93
DirectShare 32.99 36.25 57.61 50.00 47.29
30% Magnitude 33.78 33.75 46.65 50.00 51.99
LLM-Pruner 32.99 34.38 57.16 54.21 45.85
DirectShare 32.99 35.00 58.33 50.00 46.57
Table 1: Evaluation Results on Reasoning of the Memory-efficient Llama 2-7B & Llama 2-13B.
Table 2 shows the detailed results of  DirectShare applied in Llama 2 model family on these benchmarks. We provide a more comprehensive evaluation on Baichuan 2 models in Appendix B.2. Compared to reasoning tasks, our experimental results unveil a notable performance decrease of approximately 30% in large-scale reading comprehension datasets when applying DirectShare to Llama 2-7B model. Beyond this, we discover that on content summary and analysis tasks, DirectShare manages to retain 94.23% of the performance exhibited by the base model. The evaluation results of Llama 2-13B align with those of Llama 2-7B and we find the accuracy gap is larger as model size increases. This trend also exists in Magnitude Pruning and LLM-Pruner, even the performance drop is larger: LLM-Pruner drops $\approx$ 3 points more than ours on average while Magnitude Pruning is outperformed by ours by a large margin.
To mitigate this degradation, some post-training pruning methods like SparseGPT (Frantar and Alistarh, 2023) preserves accuracy via the weight update procedure. Similarly, LLM-Pruner uses the low-rank approximation (LoRA,  Hu et al., 2021) to post-train the pruned model. Motivated by this, our PostShare proves to be beneficial, substantially improving 17.80% accuracy, albeit at a certain amount of time cost. For more details refer to Section 5.2.2. However, this does not diminish the significance of our DirectShare. The absence of post-training allows us to better understand the feasibility of head-wise weight sharing for LLMs.
Ratio Method RACE- middle RACE- high OBQA CSL TNEWS
0% Llama 2-7B 33.15 35.51 31.80 55.62 20.22
10% Magnitude 25.42 26.47 28.20 49.38 14.85
LLM-Pruner 28.20 30.73 27.20 53.12 19.76
DirectShare 28.34 28.96 28.20 54.37 20.86
30% Magnitude 21.80 21.53 25.00 45.62 7.01
LLM-Pruner 21.52 22.21 26.80 50.00 10.20
DirectShare 21.45 21.53 26.00 51.25 20.22
0% Llama 2-13B 60.24 58.03 42.40 58.75 22.13
10% Magnitude 22.42 21.78 27.40 51.25 15.39
LLM-Pruner 51.46 50.80 47.00 56.25 20.95
DirectShare 54.04 55.63 39.40 56.88 17.94
30% Magnitude 21.80 22.01 28.80 46.25 4.19
LLM-Pruner 23.96 25.33 26.40 53.75 16.76
DirectShare 26.53 27.53 27.40 59.38 16.12
Table 2: NLU Abilities of the Memory-efficient Models.
@@
Table 3: Performance of Memory-efficient Llama 2-7B via PostShare. See Appendix C for results on Llama 2-13B.
Knowledge-related Tasks. We perform evaluations regarding knowledge on various datasets: WinoGrande (Levesque et al., 2012) about language, BoolQ (Clark et al., 2019) testing knowledge question answering, C-Eval (Huang et al., 2023) and MMLU (Hendrycks et al., 2021) standing for two comprehensive examination benchmarks. Table 4 summarizes the mean accuracies on those tasks after DirectShare applied to Llama 2 models. See Appendix B.3 for the results based on Baichuan 2 models.
As depicted in Table 4, DirectShare takes a clear advantage over other approaches in the field of examination. Our chosen C-Eval and MMLU span diverse disciplines to test both world knowledge and problem solving ability exclusively in a Chinese and English context, respectively. To make this more concrete, Figure 4 vividly contrasts the performance across different subjects based on Llama 2-7B on C-Eval and MMLU. But we have to admit directly do weight sharing across attention heads results in a obvious decline in knowledge-related abilities, which can be solved in PostShare.
Ratio Method WinoGrande BoolQ C-Eval MMLU
0% Llama 2-7B 54.04 70.67 32.20 46.69
10% Magnitude 51.58 60.80 22.16 28.20
LLM-Pruner 52.98 66.09 22.31 38.11
DirectShare 52.63 67.74 28.75 43.43
30% Magnitude 50.88 44.59 24.38 23.15
LLM-Pruner 50.88 54.77 22.82 25.16
DirectShare 50.18 54.43 26.24 26.53
0% Llama 2-13B 55.44 71.50 40.17 55.81
10% Magnitude 49.82 62.32 22.52 27.54
LLM-Pruner 55.44 68.07 30.25 51.45
DirectShare 54.39 69.45 37.17 52.81
30% Magnitude 49.12 56.45 23.99 22.86
LLM-Pruner 51.58 63.21 22.17 27.22
DirectShare 50.18 59.36 22.30 30.79
Table 4: Results on Knowledge-related Tasks of the Memory-efficient Models.
Figure 4: Performance across Different Subjects based on Llama 2-7B on C-Eval and MMLU.

### v2

@@ Abstract
Large Language Models (LLMs) suffer from huge number of parameters, which restricts their deployment on edge devices. Weight sharing is one promising solution that encourages weight reuse, effectively reducing memory usage with less performance drop. However, current weight sharing techniques primarily focus on small-scale models like BERT and employ coarse-grained sharing rules, e.g., layer-wise. This becomes limiting given the prevalence of LLMs and sharing an entire layer or block obviously diminishes the flexibility of weight sharing. In this paper, we present a perspective on head-wise shareable attention for large language models. We further propose two memory-efficient methods that share parameters across attention heads, with a specific focus on LLMs. Both of them use the same dynamic strategy to select the shared weight matrices. The first method directly reuses the pre-trained weights without retraining, denoted as DirectShare. The second method first post-trains with constraint on weight matrix similarity and then shares, denoted as PostShare. Experimental results reveal our head-wise shared models still maintain satisfactory capabilities, demonstrating the feasibility of fine-grained weight sharing applied to LLMs [^1] https://github.com/zouyingcao/DirectShare.
@@ 1 Introduction
Large Language Models (LLMs) have achieved breakthrough performance in a variety of natural language processing tasks  (Wei et al., 2022; Bubeck et al., 2023; Zhao et al., 2023). However, such remarkable capability typically comes at the cost of a substantial increase in the model size  (Kaplan et al., 2020). Thus, LLMs with billions of parameters  (Brown et al., 2020; Touvron et al., 2023) are more resource-hungry despite a wide margin of superiority over small-scale models (Devlin et al., 2019; Liu et al., 2019). This can also pose challenges for deployment on low-capability devices due to limited storage and GPU memory.
To address the high memory requirements of models, weight sharing  (Takase and Kiyono, 2021a; Liu et al., 2023) aims to reuse the same parameters to achieve memory- and storage-efficiency while preserving model performance. For small-scale models, e.g., BERT, it is known that several techniques  (Lan et al., 2020; Liu et al., 2023) are proposed to explore across-layer parameter sharing. While,  Zhang et al. (2022) demonstrate identical weights across different layers are the main cause of training instability and performance degradation. Moreover, the effectiveness of similar techniques at the scale of LLMs remains uncertain.
Thus, we strive to solve this central question: Can we design one fine-grained weight sharing strategy that can smoothly apply to large language models? For an effective memory-efficient weight sharing method tailored to LLMs, two key challenges must be tackled: a) the choice of shared modules whose weights are reused; b) the trade-off between reducing memory requirements and preserving diverse capabilities.
@@ 2.1 Memory-efficient Approaches for LLMs
With the growing size of language models, several memory-efficient techniques are proposed to solve. One line to reducing the memory consumption involves network compression, like quantization  (Bai et al., 2021; Tao et al., 2022), pruning  (Yang et al., 2022; Tao et al., 2023; Yang et al., 2024) and knowledge distillation (Wu et al., 2023; Tan et al., 2023). However, when applied to LLMs, many approaches have become infeasible  (Frantar and Alistarh, 2023). To recover performance, they require extensive post-training of the model  (Dettmers et al., 2023; Sun et al., 2023).
@@ 2.2 Weight Sharing
Weight sharing is a widely used technique  (Lan et al., 2020; Liu et al., 2023; Lv et al., 2023; Xu and McAuley, 2023) that aims to improve parameter efficiency and reduce inference memory requirements. Weight sharing enables model compression by eliminating redundant parameters and decouples computation and parameters by reusing the same parameters for multiple computations.
Task-oriented Weight Sharing. One of the prevalent tasks using weight sharing mechanisms is nerual machine translation (NMT). Tied Transformer  (Xia et al., 2019) considers model-level sharing and shares the weights of the encoder and decoder of an NMT model.  Dabre and Fujita (2019) proposes a method, which shares the weights across all Transformer layers and keeps performance in NMT. Besides,  Chi et al. (2021) bring the idea of ALBERT  (Lan et al., 2020) to the speech recognition task.
@@ 3.1 Attention Map Similarity: From Layer-wise to Head-wise
Prior researches (Xiao et al., 2019; Ying et al., 2021; Bhojanapalli et al., 2021) demonstrate the effectiveness of attention map reuse due to the high similarity of attention scores between different layers (especially for adjacency layers). Motivated by this, we delve into attention map similarity, specifically transitioning from layer-wise to head-wise analysis. To measure the evolution of the attention maps over layers and heads, we use the cosine similarity $\mathcal{S}_{cos}$. When $\mathcal{S}_{cos}$ equals one, it means that the attention maps are perfectly similar. Considering two specific self-attention layers, the cosine similarity is calculated as follows:
@@
We visualize the layer-wise and head-wise attention map similarity across three task-specific datasets: WMT14 (En-Fr) (Callison-Burch et al., 2009), CommonsenceQA (Talmor et al., 2019) and WSC (Levesque et al., 2012). As shown in Fig. 1(a) and (b), the degrees of similarity in attention scores computed in different layers and heads present a certain level of consistency across different tasks. In addition, we find that the cosine similarity values for pairs with high similarity are higher among different heads compared to layers. Specifically, the most similar self-attention layers reach a cosine similarity value of approximately 0.90, while in the case of head-wise comparisons, several pairs have a remarkable similarity of nearly 0.99.
One observation is that as the number of parameters increases, modules with high similarity exhibit variations, particularly in the fine-grained (e.g., head-wise) comparisons within large-scale pre-trained language models. Existing approaches employ "learning to share" techniques to dynamically adjust the sharing strategy (Xiao et al., 2019) or use a uniform sharing strategy but train the modified model from scratch (Ying et al., 2021; Shim et al., 2023). However, such strategies pay little attention to reusing attention map among heads and incur high computational costs for LLMs.
@@ 3.2 From Attention Map Similarity to Weight Matrix Similarity
Combined with the analysis of weight matrix similarity, we have made a key discovery: given a pre-trained LLM, by concatenating the weight matrix $W^{q}$ and $W^{k}$ for each head to measure the cosine similarity, the most similar weight matrix corresponds to the overlapping modules with highly similar attention maps observed across different datasets. As illustrated in Fig. 1(b) and (c), deep green circles mark the connection between attention map similarity and weight similarity (more analysis in Appendix B).
@@ 4.1 Head-wise Weight Sharing Strategy
To investigate the strategy of weight sharing applied to all the above three weight matrices across heads for LLMs, we perform preliminary experiments in the choice of head-wise match functions Match($\cdot$,$\cdot$). For the match functions, inputs are the weight matrices of head $\bm{i,j}$ and outputs are called matching scores $\bm{m}$. The higher the score, the more likely it is to share parameters across the heads.
$\bm{m_{i,j}^{*}}=\textbf{Match(}W_{i}^{*},W_{j}^{*}\textbf{)},*\subseteq\{q,k,v\}$ (2)
@@
Besides, we try another five match functions to compare: (1) Only $W^{q}_{i}$ used to measure the cosine similarity, i.e., $\bm{m_{i,j}^{*}}=\mathcal{S}_{cos}(W^{q}_{i},W^{q}_{j})$; (2) Only $W^{k}_{i}$ used to measure the cosine similarity, i.e., $\bm{m_{i,j}^{*}}=\mathcal{S}_{cos}(W^{k}_{i},W^{k}_{j})$; (3) Only $W^{v}_{i}$ used to measure the cosine similarity, i.e., $\bm{m_{i,j}^{*}}=\mathcal{S}_{cos}(W^{v}_{i},W^{v}_{j})$; (4) Concatenate all the three matrices and then calculate the cosine similarity, i.e., $\bm{m_{i,j}^{*}}=\mathcal{S}_{cos}(W^{q}_{i}||W^{k}_{i}||W^{v}_{i},W^{q}_{j}||W^{k}_{j}||W^{v}_{j})$; (5) Separately use $W^{q}_{i},W^{k}_{i},W^{v}_{i}$ to measure the cosine similarity and do weight sharing respectively, i.e., $\bm{m_{i,j}^{*}}=\mathcal{S}_{cos}(W^{*}_{i},W^{*}_{j})$ and again $*\in\{q,k,v\}$.
@@ 4.2 DirectShare
In practice, we traverse all head pairs to compute matching scores on Equation 3 and for each head, select the one with the highest score to match. When candidate shareable head pairs prepared, we select the top-N pairs in descending order according to the desired sharing ratio $\alpha$. Finally, we can share the weight matrices together between each selected attention head pairs. A detailed algorithm for our DirectShare is presented in Algorithm 1 and Appendix A.
Input: Sharing ratio $\alpha$, Original LLM $\mathcal{M}$,
Number of layers $\mathcal{L}$,
Number of heads per MHA block $\mathcal{H}$
Output: The LLM $\mathcal{M^{*}}$ after weight sharing
1 Initialize candidate buffer $\mathcal{D}_{\tau}$;
2 for $layer_{i}\leftarrow 2$ to $\mathcal{L}$ do
3      for $i\leftarrow 1$ to $\mathcal{H}$ do
4           $index_{i}\leftarrow$ $(layer_{i},i)$;
5           $index_{m}$leftarrow None;
6           $s_{m}$leftarrow -$1$;
@@
8                for $j\leftarrow 1$ to $\mathcal{H}$ do
9                     $index_{m}\leftarrow(layer_{j},j)$;
10                     Compute $\bm{\mathcal{S}_{cos}}$ using Eq. 3;
11                     if $\bm{\mathcal{S}_{cos}}>s_{m}$ then
12                          $s_{m}\leftarrow\bm{\mathcal{S}_{cos}}$
13
@@ 12                          $s_{m}\leftarrow\bm{\mathcal{S}_{cos}}$
15          Store candidate shareable head pair $<index_{i},index_{m},s_{m}>$ in $\mathcal{D}_{\tau}$;
16
17
18Sort $\mathcal{D}_{\tau}$ by descending matching scores $s_{m}$;
19 $\mathcal{N}$leftarrow Top_N($\mathcal{D}_{\tau}$, $\mathcal{L}$, $\mathcal{H}$, $\alpha$);
$\mathcal{M^{*}}$leftarrow $\mathcal{M}$, $\mathcal{N}$).
@@ 4.3 PostShare
$\mathcal{L}_{w}=\frac{1}{|\mathcal{N}|}\sum_{(i,j)\in\mathcal{N}}\left(\sum_{*\in\{q,k,v\}}\left\|W_{i}^{*}-W_{j}^{*}\right\|_{2}\right)$ (4)
where $\mathcal{N}$ is the set of selected attention head pairs for sharing. With this regularization weight loss, the proposed PostShare learn model weights by minimizing the following combined loss function:
$\mathcal{L}=\mathcal{L}_{post-training}+\gamma\times\mathcal{L}_{w}$ (5)
@@ 5 Experiments
Benchmark Type Reasoning NLU Knowledge
Ratio Method CMNLI OCNLI AX-b AX-g RTE RACE- middle RACE- high OBQA   CSL TNEWS Wino- Grande BoolQ C-Eval MMLU
0% Llama2-7B 32.98 33.12 53.53 55.34 49.82 33.15 35.51 31.80 55.62 20.22 54.04 70.67 32.20 46.69
10% Magnitude 32.99 30.63 56.70 49.44 47.29 25.42 26.47 28.20 49.38 14.85 51.58 60.80 22.16 28.20
LLM-Pruner 32.99 33.75 57.61 50.00 48.38 28.20 30.73 27.20 53.12 19.76 52.98 66.09 22.31 38.11
DirectShare 33.00 32.50 54.17 51.97 50.90 28.34 28.96 28.20 54.37 20.86 52.63 67.74 28.75 43.43
30% Magnitude 33.16 35.00 54.71 50.56 46.93 21.80 21.53 25.00 45.62 7.01 50.88 44.59 24.38 23.15
LLM-Pruner 32.99 31.25 56.34 52.53 48.74 21.52 22.21 26.80 50.00 10.20 50.88 54.77 22.82 25.16
DirectShare 33.33 32.50 57.07 51.69 49.10 21.45 21.53 26.00 51.25 20.22 50.18 54.43 26.24 26.53
0% Llama2-13B 32.99 35.00 58.81 50.56 47.29 60.24 58.03 42.40 58.75 22.13 55.44 71.50 40.17 55.81
10% Magnitude 32.82 33.12 51.99 50.56 48.38 22.42 21.78 27.40 51.25 15.39 49.82 62.32 22.52 27.54
LLM-Pruner 32.99 36.25 58.70 50.00 46.93 51.46 50.80 47.00 56.25 20.95 55.44 68.07 30.25 51.45
DirectShare 32.99 36.25 57.61 50.00 47.29 54.04 55.63 39.40 56.88 17.94 54.39 69.45 37.17 52.81
30% Magnitude 33.78 33.75 46.65 50.00 51.99 21.80 22.01 28.80 46.25 4.19 49.12 56.45 23.99 22.86
LLM-Pruner 32.99 34.38 57.16 54.21 45.85 23.96 25.33 26.40 53.75 16.76 51.58 63.21 22.17 27.22
DirectShare 32.99 35.00 58.33 50.00 46.57 26.53 27.53 27.40 59.38 16.12 50.18 59.36 22.30 30.79
Table 1: Evaluation results of DirectShare based on the Llama2-7B and Llama2-13B models. Bold and underline indicate the best and the second best results.
@@ 5.1 Experimental Settings
Models. We evaluate DirectShare and PostShare on two open-source LLMs: Llama2 (Touvron et al., 2023) and Baichuan2 (Baichuan, 2023) with 7B and 13B parameters. In PostShare, we use English Wikipedia (Foundation, ) to post-train the backbone models for weight sharing.
@@
Baselines. Since existing weight sharing techniques do not support LLMs, we compare DirectShare against Magnitude Pruning (Zhu and Gupta, 2017) and LLM-Pruner (Ma et al., 2023), two influential works for model pruning. Certainly, they are not directly comparable. To ensure fairness in the experiments, both of them only prune the multi-head attention module and thus compare when the same number of parameters is reduced. See Appendix C for additional information.
@@ 5.2.1 Evaluation on DirectShare
Table 1 shows the overall performance of DirectShare based on Llama2 models. Benchmarks are classified into three categories: reasoning, natural language understanding (NLU) and knowledge-related. The corresponding results for Baichuan2 models can be found in Appendix D.
Logical and Common Sense Reasoning. In the domain of reasoning, when applying a 30% parameter sharing to Llama2-7B, our DirectShare can still maintain an average performance of 99.51% across the five benchmarks, compared to the base model. With the same setting, the shared Llama2-13B retains 99.21% performance. This suggests our finding of head-wise shareable attention for LLMs indeed can work without significant performance degradation in reasoning tasks.
@@
Natural Language Understanding (NLU). Compared to reasoning tasks, our experimental results unveil a notable performance decrease of approximately 30% in large-scale reading comprehension datasets when applying DirectShare to Llama2-7B model. Beyond this, we discover that on content summary and analysis tasks, DirectShare manages to retain 94.23% of the performance exhibited by the base model. The evaluation results of Llama2-13B align with those of Llama2-7B and we find the accuracy gap is larger as model size increases. This trend also exists in Magnitude Pruning and LLM-Pruner, even the performance drop is larger: LLM-Pruner drops $\approx$ 3 points more than ours on average while Magnitude Pruning is outperformed by ours by a large margin.
To mitigate this degradation, some post-training pruning methods like SparseGPT (Frantar and Alistarh, 2023) preserves accuracy via the weight update procedure. Similarly, LLM-Pruner uses the low-rank approximation (LoRA,  Hu et al., 2022) to post-train the pruned model. Motivated by this, our PostShare proves to be beneficial, substantially improving 17.80% accuracy, albeit at a certain amount of time cost. For more details refer to Section 5.2.2. However, this does not diminish the significance of our DirectShare. The absence of post-training allows us to better understand the feasibility of head-wise weight sharing for LLMs.
@@
Table 2: Overall Performance of PostShare based on Llama2-7B model. See Appendix E for results on Llama2-13B.
Knowledge-related Tasks. As depicted in Table 1, DirectShare takes a clear advantage over other approaches in the field of examination. Our chosen C-Eval and MMLU span diverse disciplines to test both world knowledge and problem solving ability exclusively in a Chinese and English context, respectively. To make this more concrete, Figure 4 vividly contrasts the performance across different subjects based on Llama2-7B on C-Eval and MMLU. But we have to admit directly do weight sharing across attention heads results in a obvious decline in knowledge-related abilities, which can be solved in PostShare.
Figure 4: Performance of DirectShare across different subjects based on Llama2-7B on C-Eval and MMLU.